I have a question about how you simplify fractions with variables. Here is a problem with its solution. Why is the answer not simplified into a mixed number?
3 over 10q plus 9 over 10 q equals 6 over 5 q?


Answer by stanbon(48546)

3/10q + 9/10q = 6/5q
-----------
The lcm of the denominators is 10q
Multiply thru by 10q to get:
3 + 9 = 12
12= 12
-------------
This means the eqation is true for all values of q.
However q cannot be zero because q is in the denominator.
-------------
cheers,
Stan H.